# What Makes a Good Speech Tokenizer for LLM-Centric Speech Generation? A Systematic Study

### v2

@@ 4.3 Speaker-Aware Speech Generation
Table 2 presents that Speaker-Aware training exhibits consistent enhancement effects across NTP and MTP architectures with decoupled tokenizer.
@@
This indicates that incorporating speaker embeddings enables the model to better capture speaker’s paralinguistic features and linguistic accuracy.
@@ 4.4 Role-Playing Knowledge QA
We evaluate Role-Playing Knowledge QA within two aspects: 1) Benchmark of the decoupled SLM against top-3 coupled/semi-decoupled SLMs (From Table 1) and the LLM backbone(Qwen2.5-0.5B-Instruct); 2) Compare the decoupled SLM against diverse SLM baselines.
@@
As shown in Table 3, SLM trained with decoupled tokenizer achieves SOTA performance across all metrics. For in-domain evaluation, it outperforms all baselines with 12.0 EM and 23.8 F1, while its 0.61 SIM score is 2.9× higher than the best coupled baseline (BigCodec). Though all models degrade on OOD queries, the decoupled model shows greater robustness, confirming that decoupling resolves target conflicts in semi-decoupled/coupled systems by jointly optimizing semantic and acoustic objectives.
@@
Comparisons with other SLM baselines also confirm that decoupled tokenizers help not only improve the speech quality but preserve LLM knowledge. Detailed results are provided in Appendix E.
@@ 4.5 Ablation Study
We explore the best number of decoupled tokenizer codebook layers and speech token organization patterns for SLM training. Results show that 3 codebook layers and Frame-Wise Interleaving organization pattern is the best-performing setting. Details are provided in Section F.2 and Section F.3.
@@ 4.6 Effect of MTP on Cross-modal Alignment
Table 1 and Table 2 shows that higher speech token compression in the MTP architecture consistently improves performance, especially in WER. To investigate the hypothesis that higher compression improves modal alignment, we conduct detailed quantitative and qualitative analysis.
@@
As shown in Table 4, token compression improves cross-modal alignment and maintains intra-modal integrity, with MTP models outperforming NTP baselines. First, when compression increases from 3× to 12×, intra-modal similarities (TTSim, SSSim) remain high in all settings, while cross-modal similarity (STSim) increases steadily from 0.967 to 0.986. Second, the cross-model distance (STDist) decreases significantly from 0.244 to 0.158 , while intra-modal distances (TTDist, SSDist) stay stable. These trends align with WER improvements, suggesting that stronger compression encourages more isomorphic and effective speech-text representations. In addition, compared to NTP, all MTP settings show better cross-modal alignment. This demonstrates that MTP not only maintains intra-modal coherence but also enhances cross-modal correspondence under token compression.
Moreover, we use Riemannian distance (2009) to quantify speech-text alignment, capturing the intrinsic geometry of the embedding space by modeling it as a curved manifold. The distance between the dominant speech hidden states and text hidden states is computed as:
where is the -th positive real eigenvalue of , and are the centroids of speech and text embeddings respectively. Here can be calculated from the SVD of the right singular matrices : With higher compression, the Riemannian distance decreases (288.39 to 237.55), indicating that aggressive compression not only reduces embedding distance but also better preserves their intrinsic geometric structure.
@@ 5 Related Work
has advanced through RVQ-based architectures and quantization innovations: SoundStream (2022) introduced the RVQ-VAE framework, enhanced by Encodec (2023, LSTM temporal modeling) and SpeechTokenizer (2024, HuBERT semantic distillation). Recent ultra-low-bitrate methods include Improved RVQGAN (DAC, 2023) with quantizer dropout, WavTokenizer (2025, 0.9kbps via single-layer RVQ), BigCodec (2024, 1.04kbps via Convolution-LSTM projection), and StableCodec (2024, 0.7kbps using Transformer with FSQ (2024)).
@@
LLMs are increasingly adopted for unified speech processing through two primary paradigms. The first integrates speech encoders with LLMs for comprehension tasks like ASR (Ma et al. 2024) and spoken language understanding (Chu et al. 2023), achieving strong analytical capabilities but lacking generative functionality. The second targets speech generation via dual strategies: codec-based models (e.g., SpeechGPT, 2023) discretize speech into compressed tokens, while diffusion-enhanced approaches (e.g., CosyVoice, 2024a; 2024c) synthesize waveforms through LLM-diffusion.
@@ A.1 Compared Baselines
Encodec (Défossez et al. 2023): An RVQ-based codec designed for universal audio compression.666https://huggingface.co/facebook/encodec˙24khz
WavTokenizer (Ji et al. 2025): A single VQ codebook-based tokenizer for universal audio.777https://huggingface.co/novateur/WavTokenizer-large-speech-75token
StableCodec (Parker et al. 2024): A residual FSQ-based tokenizer for speech.888https://huggingface.co/stabilityai/stable-codec-speech-16k
BigCodec (Xin et al. 2024): A VQ-based single-stream codec for speech.999https://huggingface.co/Alethia/BigCodec/resolve/main/bigcodec.pt
SpeechTokenizer (Zhang et al. 2024): An RVQ-based codec with semantic distillation for speech.101010https://huggingface.co/fnlp/AnyGPT-speech-modules/tree/main/speechtokenizer
@@ A.2 Results of Codec Reconstruction
Table 5 shows the results of VQ reconstruction evaluation on the LibriTTS test-clean dataset. The results reveal that for acoustic quality (STOI (Taal et al. 2010), PESQ (Rix et al. 2001) and UTMOS), all tokenizers perform similar with the ground truth, sharing close metric values. But for SIM, the trend differs: the VQ performs better on SIM as its token rate increases, and the decoupled model (FACodec) surpasses most of the coupled and semi-decoupled models. This is probably because reduced token rates compromise timbre fidelity, combined with coupled modeling diminishing acoustic prominence relative to semantic components.
@@ Appendix B Speaker Embedding Extraction
In FACodec, speaker information is extracted by a timbre extractor which is composed of 4 transformer encoder blocks. The module operates on the output of the speech encoder and transforms it into a global vector , which captures the characteristics of speakers. To generate audios, we sum the representations of prosody, content, and then integrate the speaker embedding via conditional layer normalization, enabling effective conditioning on speaker timbre during generation.
@@ Appendix C Experimental Setups
For Role-Playing Knowledge QA, we leveraged the proposed Emilia-3.5 (2,495 hours) and RoleTriviaQA (138k) datasets during stage 1 and stage 2. Our proposed data details can be found in Section 3.
@@
We selected the current state-of-the-art speech tokenizer as the baseline for our model. To ensure a fair comparison, we employed the official weight files provided by the Encodec (Défossez et al. 2023), SpeechTokenizer (Zhang et al. 2024), WavTokenizer (Ji et al. 2025), StableCodec (Parker et al. 2024), BigCodec (Xin et al. 2024). Details can be found in Section A.1.
@@
Our baseline training configuration employs the AdamW optimizer ( and ) with a 5e-4 initial learning rate following a cosine decay schedule. All models undergo 90k training steps using consistent batch sizes of 16 for fair comparison. For specialized tasks like Role-Playing Knowledge QA, we implement a two-phase adaptation: 40k steps for cross-modal alignment pretraining followed by 12k steps of knowledge-role joint fine-tuning. During inference, we apply greedy decoding with a 1.2 repetition penalty to balance output quality and diversity.
@@
Figure 4 depicts the training loss of NTP TTS training with LibriTTS and 2-stage training of Role-Playing Knowledge QA task of all coupled, semi-decoupled baselines and FACodec, on which tokenizer types are annotated. It can be clearly seen that decoupled and semi-decoupled SLMs perform similarly and significantly better than coupled baselines across all tasks and stages.
@@
For Role-Playing Knowledge QA task, we report three metrics: Exact Match (EM) and F1 score for text responses, while SIM for speech responses. Details of these metrics can be found in Table 6.
@@ Appendix D Details of Evaluation Metrics
Word Error Rate (WER) measures the accuracy of the semantic information of the generated speech. We employed the ASR model Whisper-large-v3111111https://huggingface.co/openai/whisper-large-v3 to transcribe generated speech into text and then calculate the WER between the transcribed text and the text response.
UTMOS assesses the acoustic quality of the generated speech. We utilized the model provided by SaruLab121212https://huggingface.co/spaces/sarulab-speech/UTMOS-demo/tree/main to evaluate it.
@@ Appendix E RoleTriviaQA Results on SLM Baselines
Results in Table 6 show that the best baseline (SLM trained with decoupled tokenizer FACodec) in our study reaches remarkably higher EM and F1 (12.0 and 23.8) than other SLM baselines, most of which fail to generate exactly correct answers (EM=0). This result confirms the superiority of decoupled SLMs to other baselines with different architechtures.
@@ F.2 Effect of Number of Codebook Layers in SLMs
3 | CWI | 0.80 | 3.30 | 22.21 | 0.43 | ||||
3 | FWI | 0.87 | 3.93 | 6.07 | 0.49 |
As indicated in Table 8, the 3-layer setting (1 prosody layer and 2 content layers) in FACodec-NTP training exhibits superior overall performance and higher success rate compared to the full 6-layer (adding 3 acoustic layers based on the previous setting) configuration. Therefore, we chose the 3-layer setting for main NTP and MTP experiments.
@@ F.3 Effect of Speech Token Organization Pattern
Table 9 shows that FWI outperforms CWI on all metrics especially on WER (6.07 vs. 22.21), which means the CWI‐trained model aligns text and speech modalities less effectively than its FWI‐trained counterpart. Therefore, we opt FWI as the default setting for our main experiments.
@@ F.4 Effect of RoleTriviaQA data format
Table 10 compares these formats for SLMs trained with StableCodec and FACodec under both NTP and MTP settings131313Speaker embedding is unavailable for coupled tokenizers..
@@ G.2 Effect of Tokenizer-MTP Combination
Table 11 investigates the impact of different tokenizers trained on MTP based SLMs. We select semi-decoupled (SpeechTokenizer), decoupled (FACodec) tokenizers for comparsion, and choose BigCodec as our coupled tokenizer baseline, due to its 80 TPS for fair analysis and strong performance in coupled tokenizers as shown in Table 1.
Primarily, coupled models experience severe degradation as compression increases: UTMOS drops sharply from 3.97 to 1.94, and WER rises to 46.82 at 8× compression. Additionally, semi-decoupled models show some WER improvement but with a slight decline in UTMOS, even with 10x compression rate. Finally, the decoupled architecture maintains speaker similarity, and consistently reduces the WER under higher compression, with only a minor reduction in audio quality. Remarkably, even at a 12 compression, the WER is close to ground truth. Our above results indicate that: MTP architectures achieve the best performance with decoupled tokenizers, followed by semi-decoupled and coupled tokenizers.
@@ H.1 Data Construction
Table 12 demonstrates the detailed composition of RoleTriviaQA along with the selected speakers. The ID evaluation samples are collected from TriviaQA test set and OOD samples are gathered from Web Questions subset of OpenAudioBench. Notably, we added about 26,000 text QA training samples filtered from LLaVA-Instruct-150k141414https://huggingface.co/datasets/liuhaotian/LLaVA-Instruct-150K during training to preserve the original textual ability of the SLM.
@@
For each character’s speech synthesis, we select high-quality reference speeches by combining speaker embedding similarity metrics with manual filtering as an comprehensive criteria. Then we leverage the CosyVoice 2 151515https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B system to generate role-playing speech from text answers. Finally, all speeches are resampled to 16 kHz to fit the tokenization requirement.
@@ Appendix I Multi Token Prediction
Multi Token Prediction (MTP) is leveraged in LLMs to accelerate inference through three principal strategies: 1) Multi-head architectures enhance generation quality via parallel output heads (Gloeckle et al. 2024); 2) Tree-based systems like MEDUSA (Cai et al. 2024) refine candidate selection through hierarchical validation; 3) Scalable frameworks such as DeepSeek-V3 (DeepSeek-AI et al. 2024) enable efficient large-scale training. Recent adaptations extend MTP to speech processing: SLAM-Omni (Chen et al. 2024a) employs audio token grouping for computational efficiency but neglects paralinguistic modeling through semantic-focused processing, while VocalNet (Wang et al. 2025) achieves localized acceleration via DeepSeek-V3-style parallel prediction yet faces pipeline complexity from flow-matching requirements.

### v3

@@ 4.3 Speaker-Aware Speech Generation
Table 2 presents that Speaker-Aware training exhibits consistent enhancement effects across NTP and MTP architectures with a decoupled tokenizer.
@@
This indicates that incorporating speaker embeddings enables the model better to capture the speaker’s paralinguistic features and linguistic accuracy.
@@ 4.4 Role-Playing Knowledge QA
We evaluate Role-Playing Knowledge Speech QA within two aspects: 1) Benchmark of the decoupled SLM against top-3 coupled/semi-decoupled SLMs (From Table 1) and the LLM backbone (Qwen2.5-0.5B-Instruct); 2) Compare the decoupled SLM against diverse SLM baselines.
@@
Table 3 shows that SLM trained with a decoupled tokenizer achieves SOTA performance across all metrics. For ID evaluation, it outperforms all baselines with 12.0 EM and 23.8 F1, while its 0.61 SIM score is higher than the best coupled baseline (BigCodec). Though all models degrade on OOD queries, the decoupled model shows greater robustness, confirming that decoupling resolves target conflicts in semi-decoupled/coupled systems by jointly optimizing semantic and acoustic objectives.
@@
Comparisons with other SLM baselines also confirm that decoupled tokenizers help not only improve the speech quality but also preserve LLM knowledge. Detailed results are provided in Appendix E.
@@ 4.5 Ablation Study
We explore the best number of decoupled tokenizer codebook layers and speech token organization patterns for SLM training. Results show that 3 codebook layers with Frame-Wise Interleaving organization pattern is the best-performing setting. Details are provided in Section F.2 and Section F.3.
@@ 4.6 Effect of MTP on Cross-modal Alignment
Table 1 and Table 2 show that higher speech token compression in the MTP architecture consistently improves performance, especially in WER. To investigate the hypothesis that higher compression improves modal alignment, we conduct detailed quantitative and qualitative analyses.
@@
As shown in Table 4, token compression improves cross-modal alignment and maintains intra-modal integrity, with MTP models outperforming NTP baselines. First, when compression increases from to , intra-modal similarities (TTSim, SSSim) remain high in all settings, while cross-modal similarity (STSim) increases steadily from 0.967 to 0.986. Second, the cross-model distance (STDist) decreases significantly from 0.244 to 0.158 , while intra-modal distances (TTDist, SSDist) stay stable. These trends align with WER improvements, suggesting that stronger compression encourages more isomorphic and effective speech-text representations. Additionally, compared to NTP, all MTP settings exhibit better cross-modal alignment. This demonstrates that MTP not only maintains intra-modal coherence but also enhances cross-modal correspondence under token compression.
We also use Riemannian distance (2009) to quantify speech-text alignment (See Table 6 for details). Table 4 shows that with higher compression, the Riemannian distance decreases (288.39 to 237.55), indicating that aggressive compression not only reduces embedding distance but also better preserves their intrinsic geometric structure.
@@ 5 Related Work
has advanced through RVQ-based architectures and quantization innovations: SoundStream (2022) introduced the RVQ-VAE framework, enhanced by Encodec (2023, LSTM temporal modeling) and SpeechTokenizer (2024, HuBERT semantic distillation). Recent ultra-low-bitrate methods include Improved RVQGAN (DAC, 2023) with quantizer dropout, WavTokenizer (2025, 0.9kbps via single-layer RVQ), BigCodec (2024, 1.04kbps via Convolution-LSTM projection), and StableCodec (2025, 0.7kbps using Transformer with FSQ (2024)).
@@
LLMs are increasingly adopted for unified speech processing through two primary paradigms. The first integrates speech encoders with LLMs for comprehension tasks like ASR (Ma et al. 2024) and spoken language understanding (Chu et al. 2023), achieving strong analytical capabilities but lacking generative functionality. The second targets speech generation via dual strategies: codec-based models (e.g., SpeechGPT, 2023) discretize speech into compressed tokens, while diffusion-enhanced approaches (e.g., CosyVoice, 2024a; 2024a) synthesize waveforms through LLM-diffusion.
@@ 6 Conclusion
Acknowledgements
The authors wish to thank the anonymous reviewers for their helpful comments. This work was partially funded by National Natural Science Foundation of China (No.62476061, 62376061,62576106, 62506079), Shanghai Rising-Star Program (23QA1400200), and Natural Science Foundation of Shanghai (23ZR1403500). The computations in this research were performed using the CFFF platform of Fudan University.
@@ A.1 Compared Baselines
Encodec (Défossez et al. 2023): An RVQ-based codec designed for universal audio compression.555https://huggingface.co/facebook/encodec˙24khz
WavTokenizer (Ji et al. 2025): A single VQ codebook-based tokenizer for universal audio.666https://huggingface.co/novateur/WavTokenizer-large-speech-75token
StableCodec (Parker et al. 2025): A residual FSQ-based tokenizer for speech.777https://huggingface.co/stabilityai/stable-codec-speech-16k
BigCodec (Xin et al. 2024): A VQ-based single-stream codec for speech.888https://huggingface.co/Alethia/BigCodec/resolve/main/bigcodec.pt
SpeechTokenizer (Zhang et al. 2024): An RVQ-based codec with semantic distillation for speech.999https://huggingface.co/fnlp/AnyGPT-speech-modules/tree/main/speechtokenizer
@@ A.2 Results of Codec Reconstruction
Table 5 shows the results of VQ reconstruction evaluation on the LibriTTS test-clean dataset. The results reveal that for acoustic quality (STOI (Taal et al. 2010), PESQ (Rix et al. 2001) and UTMOS), all tokenizers perform similarly with the ground truth, sharing close metric values. But for SIM, the trend differs: the VQ performs better on SIM as its token rate increases, and the decoupled model (FACodec) surpasses most of the coupled and semi-decoupled models. This is probably because reduced token rates compromise timbre fidelity, combined with coupled modeling diminishing acoustic prominence relative to semantic components.
@@ Appendix B Speaker Embedding Extraction
In FACodec, speaker information is extracted by a timbre extractor, which consists of four transformer encoder blocks. The module operates on the output of the speech encoder and transforms it into a global vector , which captures the characteristics of speakers. To generate audio, we sum the representations of prosody and content, and then integrate the speaker embedding via conditional layer normalization, enabling effective conditioning on speaker timbre during generation.
@@ Appendix C Experimental Setups
For Role-Playing Knowledge QA, we utilized the proposed Emilia-3.5 (2,495 hours) and RoleTriviaQA (138k) datasets during stages 1 and 2. Our proposed data details can be found in Section 3.
@@
We selected the current state-of-the-art speech tokenizer as the baseline for our model. To ensure a fair comparison, we employed the official weight files provided by the Encodec (Défossez et al. 2023), SpeechTokenizer (Zhang et al. 2024), WavTokenizer (Ji et al. 2025), StableCodec (Parker et al. 2025), BigCodec (Xin et al. 2024). Details can be found in Section A.1.
@@
Our baseline training configuration employs the AdamW optimizer ( and ) with an initial learning rate of 5e-4 following a cosine decay schedule. All models undergo 90k training steps using consistent batch sizes of 16 for fair comparison. For specialized tasks like Role-Playing Knowledge QA, we implement a two-phase adaptation: 40k steps for cross-modal alignment pretraining followed by 12k steps of knowledge-role joint fine-tuning. During inference, we apply greedy decoding with a 1.2 repetition penalty to balance output quality and diversity.
@@
Figure 4 depicts the training loss of NTP TTS training with LibriTTS and 2-stage training of the Role-Playing Knowledge QA task of all coupled, semi-decoupled baselines and FACodec, on which tokenizer types are annotated. It can be clearly seen that decoupled and semi-decoupled SLMs perform similarly and significantly better than coupled baselines across all tasks and stages.
@@
For the Role-Playing Knowledge QA task, we report three metrics: Exact Match (EM) and F1 score for text responses, while SIM for speech responses. Details of these metrics can be found in Table 6.
@@ Appendix D Details of Evaluation Metrics
Word Error Rate (WER) measures the accuracy of the semantic information of the generated speech. We employed the ASR model Whisper-large-v3101010https://huggingface.co/openai/whisper-large-v3 to transcribe generated speech into text and then calculate the WER between the transcribed text and the text response.
UTMOS assesses the acoustic quality of the generated speech. We utilized the model provided by SaruLab111111https://huggingface.co/spaces/sarulab-speech/UTMOS-demo/tree/main to evaluate it.
@@
We also utilized Riemannian Distance to quantify speech-text alignment.
Riemann Distance captures the intrinsic geometry of the embedding space by modeling it as a curved manifold. The distance between the dominant speech hidden states and text hidden states is computed as:
where is the -th positive real eigenvalue of , and are the centroids of speech and text embeddings respectively. Here can be calculated from the SVD of the right singular matrices :
@@ Appendix E RoleTriviaQA Results on SLM Baselines
Results in Table 6 show that the best baseline (SLM trained with decoupled tokenizer FACodec) in our study reaches remarkably higher EM and F1 (12.0 and 23.8) than other SLM baselines, most of which fail to generate exactly correct answers (EM=0). This result confirms the superiority of decoupled SLMs to other baselines with different architectures.
@@ F.2 Effect of Number of Codebook Layers in SLMs
| 3 | CWI | 0.80 | 3.30 | 22.21 | 0.43 | ||||
| 3 | FWI | 0.87 | 3.93 | 6.07 | 0.49 |
As indicated in Table 8, the 3-layer setting (1 prosody layer and 2 content layers) in FACodec-NTP training exhibits superior overall performance and a higher success rate compared to the full 6-layer (adding 3 acoustic layers based on the previous setting) configuration. Therefore, we chose the 3-layer setting for the main NTP and MTP experiments.
@@ F.3 Effect of Speech Token Organization Pattern
Table 9 shows that FWI outperforms CWI on all metrics, especially on WER (6.07 vs. 22.21), which means the CWI‐trained model aligns text and speech modalities less effectively than its FWI‐trained counterpart. Therefore, we opt for FWI as the default setting for our main experiments.
@@ F.4 Effect of RoleTriviaQA data format
Table 10 compares these formats for SLMs trained with StableCodec and FACodec under both NTP and MTP settings121212Speaker embedding is unavailable for coupled tokenizers..
@@ G.2 Effect of Tokenizer-MTP Combination
Table 11 investigates the impact of different tokenizers trained on MTP-based SLMs. We select semi-decoupled (SpeechTokenizer), decoupled (FACodec) tokenizers for comparison, and choose BigCodec as our coupled tokenizer baseline, due to its 80 TPS for fair analysis and strong performance in coupled tokenizers as shown in Table 1.
Primarily, coupled models experience severe degradation as compression increases: UTMOS drops sharply from 3.97 to 1.94, and WER rises to 46.82 at 8× compression. Additionally, semi-decoupled models show some WER improvement but with a slight decline in UTMOS, even with a compression rate. Finally, the decoupled architecture maintains speaker similarity and consistently reduces the WER under higher compression, with only a minor reduction in audio quality. Remarkably, even at a compression, the WER is close to the ground truth. Our above results indicate that MTP architectures achieve the best performance with decoupled tokenizers, followed by semi-decoupled and coupled tokenizers.
@@ H.1 Data Construction
Table 12 demonstrates the detailed composition of RoleTriviaQA along with the selected speakers. The ID evaluation samples are collected from the TriviaQA test set, and OOD samples are gathered from the Web Questions subset of OpenAudioBench. Notably, we added about 26,000 text QA training samples filtered from LLaVA-Instruct-150k131313https://huggingface.co/datasets/liuhaotian/LLaVA-Instruct-150K during training to preserve the original textual ability of the SLM.
@@
For each character’s speech synthesis, we select high-quality reference speeches by combining speaker embedding similarity metrics with manual filtering as a comprehensive criterion. Then we leverage the CosyVoice 2 141414https://huggingface.co/FunAudioLLM/CosyVoice2-0.5B system to generate role-playing speech from text answers. Finally, all speeches are resampled to 16 kHz to fit the tokenization requirement.
@@ Appendix I Multi Token Prediction
Multi Token Prediction (MTP) is leveraged in LLMs to accelerate inference through three principal strategies: 1) Multi-head architectures enhance generation quality via parallel output heads (Gloeckle et al. 2024); 2) Tree-based systems like MEDUSA (Cai et al. 2024) refine candidate selection through hierarchical validation; 3) Scalable frameworks such as DeepSeek-V3 (DeepSeek-AI 2025) enable efficient large-scale training. Recent adaptations extend MTP to speech processing: SLAM-Omni (Chen et al. 2024a) employs audio token grouping for computational efficiency but neglects paralinguistic modeling through semantic-focused processing, while VocalNet (Wang et al. 2025) achieves localized acceleration via DeepSeek-V3-style parallel prediction yet faces pipeline complexity from flow-matching requirements.